Solve the equation by using the square root property. Express the radical in simplest form: (m+12)^2=-9
Answer by rothauserc(4718)

(m+12)^2=-9
take square root of both sides
m+12 = 3i
note that i = square root of -1
m = 3i -12
m = 3*(i -4)